OpenStudy (anonymous):

A student who sells subscriptions for a magazine that costs $35 a year makes a commission of $5.25 on each subscription. What percent commission does the student make?

hero (hero):

So for each sell, the customer pays $35 and the student receives $5.25. So basically, out of $35, the student gets 5.25. So this can be represented by a fraction: \[\frac{5.25}{35}\]

hero (hero):

Now let's set that equal to The percentage of 100 that we need to find: \[\frac{5.25}{35} = \frac{\%}{100}\] Since the percentage is unknown, replace it with x: \[\frac{5.25}{35} = \frac{x}{100}\]

hero (hero):

Cross multiply and solve for x

OpenStudy (anonymous):

Okay thank you very much

hero (hero):

No, you can't get off that easy. You have to tell me what x equals. Then I give you a medal.

OpenStudy (anonymous):

Haha okay.

OpenStudy (anonymous):

Does x equal 15?

hero (hero):

Yes, and remember x = %, so what does that mean?

OpenStudy (anonymous):

x equals 15%

hero (hero):

Good job. By the way... In general, when trying to figure out percentages, use the following formula: \[\frac{\text{is}}{\text{of}} = \frac{\%}{100}\]

OpenStudy (anonymous):

Okay thank you for the tip it is greatly appreciated

hero (hero):

You should be able to find any percent now. What is 14% of 5?

OpenStudy (anonymous):

7%?

hero (hero):

Actually, I already gave you the percent in this problem. You have to find the number.

OpenStudy (anonymous):

0.7

hero (hero):

Yes, good, but I want to see how you set up the problem.

OpenStudy (anonymous):

x over 5 =14 over 100. Then I crossed multiplied, in which I got 100x/70. Then I divided that and so I got 0.7

hero (hero):

Yup, that is correct. YOU have figured it out.